OpenStudy (anonymous):

3/x = 2/x-2 solving for x

OpenStudy (anonymous):

Multiply everything by x.

OpenStudy (anonymous):

what about the - 2?

OpenStudy (anonymous):

It would become -2x

OpenStudy (anonymous):

do you know how to cross multiply?

OpenStudy (anonymous):

\[ \large \frac{ 3 }{ x } = \frac{ 2 }{ x-2 }\] by cross multiply we have \[\large 3(x-2)= 2x\]

OpenStudy (anonymous):

So you would have 3= 2 -2x

OpenStudy (anonymous):

\[\large 3x -6 = 2x\] \[\large x= 6\]

OpenStudy (anonymous):

Now, solve for x.

OpenStudy (anonymous):

no, you can't multiply by x to solve because \[\large \frac{ 3 }{ x } (x) = \frac{ 2 }{ x-2 } (x)\] is just \[\large 3 = \frac{ 2 x}{ x-2 } \] then you'll have to multiply x-2 over to 3 \[3(x-2)=2x\] which will generate the same answer as mine, x=6 but cross multiply saves alot of space.

OpenStudy (anonymous):

Thank You :)

OpenStudy (anonymous):

\(\color{salmon}{\large\heartsuit}\large\text{You're Most Welcome! }\color{salmon}\heartsuit\) \(~~~~~~~~~~~~~~~\color{green}{\large\ddot\smile}\color{blue}{\large\ddot\smile}\color{salmon}{\large\ddot\smile}\color{blue}{\large\ddot\smile}\color{salmon}{\large\ddot\smile}\) \(~~~~~~~~~~~~~~~~~~~~\color{salmon}{\large\bigstar}\color{salmon}{\large\bigstar}\color{salmon}{\large\bigstar} \)

OpenStudy (anonymous):

and do you still need help with the question from message?